# StreamSplat: Towards Online Dynamic 3D Reconstruction from Uncalibrated Video Streams

### v1

@@ Abstract
Real-time reconstruction of dynamic 3D scenes from uncalibrated video streams is crucial for numerous real-world applications. However, existing methods struggle to jointly address three key challenges: 1) processing uncalibrated inputs in real time, 2) accurately modeling dynamic scene evolution, and 3) maintaining long-term stability and computational efficiency. To this end, we introduce StreamSplat, the first fully feed-forward framework that transforms uncalibrated video streams of arbitrary length into dynamic 3D Gaussian Splatting (3DGS) representations in an online manner, capable of recovering scene dynamics from temporally local observations. We propose two key technical innovations: a probabilistic sampling mechanism in the static encoder for 3DGS position prediction, and a bidirectional deformation field in the dynamic decoder that enables robust and efficient dynamic modeling. Extensive experiments on static and dynamic benchmarks demonstrate that StreamSplat consistently outperforms prior works in both reconstruction quality and dynamic scene modeling, while uniquely supporting online reconstruction of arbitrarily long video streams. Code and models are available at https://github.com/nickwzk/StreamSplat.
Figure 1: Given an uncalibrated video stream, our StreamSplat performs instant reconstruction of dynamic 3D Gaussian scene in an online manner, enabling video reconstruction and interpolation, depth estimation, and novel view synthesis.
@@ 1 Introduction
Real-time reconstruction of dynamic 3D scenes from video streams is crucial for numerous applications, e.g., robotics [1, 2], augmented/virtual reality (AR/VR) [3, 4], and autonomous driving [5]. AR/VR systems rely on accurate, continuously updated 3D models to deliver immersive user experiences, while robots and autonomous vehicles require real-time representations of dynamic environments for safe navigation and responsive interactions. These applications demand robust online reconstruction systems capable of accurately capturing evolving scene geometry and appearance.
Despite growing research interest, real-time dynamic 3D reconstruction from uncalibrated video remains largely unsolved. Although SLAM-based [6, 7, 8, 9] and scene-coordinate-based methods [10, 11, 12, 13] could effectively estimate camera poses and reconstruct static scenes, their extension to dynamic scenes is often hindered by the challenge of disentangling camera and object motion from uncalibrated inputs. As a result, they often require computationally intensive post-optimization of camera poses or scene representations [14, 6], limiting their real-time applicability. Meanwhile, feed-forward methods [15, 16, 17] are promising for static scene reconstruction from uncalibrated inputs. Recent work [18] has begun to explore their potential for dynamic scene reconstruction. However, such methods typically treat dynamic scenes as a sequence of static ones, ignoring temporal coherence and failing to explicitly model scene dynamics. Moreover, many require access to entire video sequences during inference [18, 19], making them unsuitable for online streaming scenarios.
Building such an online 3D reconstruction system for dynamic scenes introduces three fundamental challenges: (1) Real-Time Processing. Latency-sensitive tasks [1, 5] require immediate 3D reconstruction. However, traditional methods often rely on offline calibration such as Structure-from-Motion (SfM) [20] or iterative optimization of scene representations [21, 22], making them unsuitable for real-time processing of uncalibrated video streams [23, 24, 25]; (2) Dynamic Motion Modeling. In uncalibrated videos, camera and object motions are almost always entangled, and dynamic scenes exhibit complex, non-rigid motions that must be inferred from video frames [26, 27, 28]; (3) Long-Term Stability. Online systems must process arbitrarily long video streams while preventing error accumulation and maintaining efficiency in both memory and computation [17].
To address these challenges, we propose StreamSplat, the first fully feed-forward framework for online dynamic 3D Gaussian Splatting (3DGS) [21, 29] reconstruction from uncalibrated video streams. Inspired by the recent success of feed-forward Gaussian Splatting [15], we leverage pixel-aligned 3D Gaussians in an orthographic canonical space [26, 30] to support online, dynamic, and uncalibrated reconstruction. We formulate online video reconstruction as a sequential, frame-by-frame process. First, we encode each incoming frame into a set of pixel-aligned static 3D Gaussians in a probabilistic way. This formulation is specifically designed for 3DGS position prediction (Section 3.1), alleviating positional local minima caused by random initialization and better capturing uncertainty.
To model dynamic motion, we propose a novel bidirectional deformation field coupled with an adaptive Gaussian fusion mechanism. These components enable smooth transitions and seamless fusion of Gaussians across frames by softly matching spatially and temporally coherent structures, while dynamically integrating newly observed content (Section 3.2). The bidirectional formulation ensures that dynamic Gaussians remain structured and pixel-aligned at key frames, improving robustness and mitigating the error accumulation common in long-term online reconstruction [17]. We evaluate StreamSplat on both static (CO3Dv2 [31], RealEstate10K [32]) and dynamic (DAVIS [33], YouTube-VOS [34]) benchmarks. Our method consistently outperforms prior works in both reconstruction quality and dynamic scene modeling, while uniquely supporting fully online reconstruction of arbitrarily long video streams (Section 4). In summary, our key contributions are:
@@
  • We propose StreamSplat, the first fully feed-forward framework for online dynamic 3DGS reconstruction from uncalibrated video streams.
@@
  • We introduce two key technical innovations: a probabilistic sampling mechanism in the static encoder for 3DGS position prediction, and a bidirectional deformation field in the dynamic decoder that enables robust and efficient dynamic modeling.
  • StreamSplat achieves state-of-the-art reconstruction quality and dynamic scene modeling across multiple static and dynamic benchmarks, while uniquely supporting online reconstruction of arbitrarily long uncalibrated video streams.
@@ 2 Related Works
2.1 Neural Representations for Dynamic Scenes
Dynamic 3D reconstruction from monocular videos is critical for many real-world applications. Early works leveraged implicit neural representations and modeled dynamic scenes using coordinate-based multilayer perceptrons (MLPs) optimized through per-image optimization [35, 36, 37, 38]. Later works extended by introducing learnable time-conditioned deformation fields with canonical representations [39, 40, 41, 42]. However, these methods typically fail to capture large and complex motions due to lack of explicit 3D structure and motions. To address this, recent work shifts to explicit 3D primitives, notably dynamic Gaussian splatting [43, 44, 45, 46, 30, 47, 48], which leverages persistent and deformable Gaussians to efficiently represent dynamic scenes for real-time rendering. However, these methods still rely on calibrated camera poses and extensive per-scene optimization, making them unsuitable for fully uncalibrated and real-time scenarios.
2.2 Feed-Forward Dynamic 3D Reconstruction
By directly predicting 3D scene representations via neural networks, feed-forward methods have emerged as promising alternatives to optimization-based approaches. For example, based on DUSt3R [11], several approaches have shown the ability to estimate camera parameters and 3D point clouds from uncalibrated inputs [13, 49, 50, 12, 51, 52, 53]. MonST3R [14] extends this paradigm to dynamic scenes by producing temporally coherent point clouds and camera poses across frames. Such methods typically require post-optimization on camera poses or scene geometry to maintain consistency in dynamic scenarios. Moreover, scene-coordinate-based point clouds are inherently difficult to perform deformations, limiting their effectiveness in dynamic scene modeling [43, 45].
Another research line explores fully feed-forward 3D Gaussian Splatting (3DGS) methods for 3D reconstruction [54, 55, 56, 57, 58, 59, 60, 61, 62]. Prior works achieve fast static scene reconstruction from images pairs. Specifically, pixelSplat [15] employs multi-view transformers to regress 3D Gaussians from calibrated image pairs, while NoPoSplat [16] extends this to uncalibrated pairs, and StreamGS [17] achieves online static reconstruction from image streams. Recent works such as BTimer [18] further adapt these approaches to dynamic scene reconstruction by stacking multiple temporal-consistent static 3DGS from calibrated video frames. Despite their efficiency, these methods either primarily address static scenes or approximate dynamics by stacking static representations, which fail to explicitly model scene dynamics. Moreover, many require access to entire video sequences during inference [18, 19], making them unsuitable for online streaming scenarios.
@@ 3 Method
Figure 2: Overview of the StreamSplat framework. Given a pair of frames, we first encode them using the Static Encoder to produce canonical 3D Gaussians (Section 3.1), and then pass the 3DGS Embeddings to the Dynamic Decoder to predict the deformation field (Section 3.2). The resulting dynamic 3D Gaussians can be rendered at arbitrary time to produce RGB images and depth maps.
In this section, we present StreamSplat, a framework designed to instantly transforms uncalibrated online video streams into dynamic 3D Gaussian Splatting (3DGS) representations capable of capturing scene dynamics. Figure 2 provides an overview. We first encode the current frame into static 3D Gaussians in a canonical space (Section 3.1), then predict a bidirectional deformation field between Gaussians encoded from current frame and propagated from the previous frame, and finally fuse them into a unified dynamic representation (Section 3.2). This representation supports rendering at arbitrary times and viewpoints, thereby effectively recovering scene dynamics (Section 3.3).
@@ 3.1 Probabilistic 3D Gaussian Encoding
Canonical 3D Space. Following [26, 30, 19], we define the canonical space using a shared orthographic camera coordinate system. Details are provided in Appendix A. We argue that this choice offers two key advantages. First, under orthographic projection, scene dynamics that typically entangle camera and object motion manifest as a shared global shift, which can be naturally absorbed by Gaussian motions and handled by our dynamic prediction module in an unified manner (Section 3.2). This simplification allows us to bypass errors from inaccurate camera estimation. Second, orthographic projection simplifies the camera model while still preserving the spatio-temporal structure of 3D motion [26]. Together, this orthographic canonical space provides a simplified, yet effective foundation for predicting dynamic 3D representations directly from uncalibrated videos.
@@
Structured Static 3D Gaussian Encoding. To overcome the inherent depth ambiguity under orthographic projection [63] and unstructured nature of 3DGS [64], we incorporate a pretrained depth estimator [65] to generate per-frame pseudo-depth maps as ground truth and predict 3D Gaussian positions in a pixel-aligned manner [15] to maintain spatial correspondence with input frames.
@@
As shown in Figure 2, given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we first obtain the corresponding depth map and concatenate it with $I$ to form a 4-channel RGBD input. This is processed by a non-overlapping 2D convolution to extract $8\times 8$ patches, which are then fed into a Transformer encoder with self-attention and MLP blocks [66, 67], producing 3DGS Embeddings $\mathbf{h}\in\mathbb{R}^{N\times D}$, where $N=HW/64$. These embeddings are upsampled via a lightweight block composed of linear layers and window-based attention [68, 69], yielding static Gaussian tokens $\hat{\mathbf{E}}\in\mathbb{R}^{16N\times D/4}$. Each token $\hat{\mathbf{E}}_{i}$ is decoded by linear heads into 3DGS parameters: position offset $\boldsymbol{o}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{R}_{i}\in\mathbb{R}^{4}$, scale $\mathbf{S}_{i}\in\mathbb{R}^{3}$, opacity $\boldsymbol{\alpha}_{i}\in\mathbb{R}$, and color $\boldsymbol{c}_{i}\in\mathbb{R}^{3}$. The final 3D position $\boldsymbol{\mu}_{i}$ is computed via pixel-aligned prediction: $\boldsymbol{\mu}_{i}=\left(u+\boldsymbol{o}_{i,0},v+\boldsymbol{o}_{i,1},1/\boldsymbol{o}_{i,2}\right)$, where $(u,v)$ is the pixel coordinate of the $i$-th token. The first two offset components $\boldsymbol{o}_{i,0}$, $\boldsymbol{o}_{i,1}$ specify the local offset within the $2\times 2$ image patch, while $\boldsymbol{o}_{i,2}$ denotes the inverse depth following [65] for a better depth estimation near the camera.
@@
Probabilistic Position Sampling. 3D Gaussian Splatting is sensitive to position initialization [21] and prone to local minima [15], especially in feed-forward models [15, 58] that make predictions in a single forward pass without iterative refinement. Inspired by [15], we predict a truncated normal distribution for each 3D offset $\boldsymbol{o}$ rather than regressing it directly: $\boldsymbol{o}\sim\mathcal{N}_{[-1,1]}(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p})$, where $\boldsymbol{\mu}_{p}$ and $\boldsymbol{\Sigma}_{p}$ are the predicted mean and covariance. As shown in Section 4.4, this strategy promotes spatial exploration during early training and stabilizes convergence toward optimal positions.
@@ 3.2 Dynamic Deformation Prediction
Bidirectional Deformation Field. Online dynamic scene reconstruction involves complex motion, including large deformations and emerging objects/surfaces [17]. To handle this, we propose a bidirectional deformation field that jointly models forward and backward temporal motion between consecutive frames. Specifically, our model estimates: (i) a forward deformation field that transforms the current-frame Gaussians $\mathcal{G}_{t}$ toward the next frame, and (ii) a backward deformation field that transforms next-frame Gaussians $\mathcal{G}_{t+1}$ back to the current frame.
Each deformation field consists of a 3D velocity vector $\mathbf{v}\in[-1,1]^{3}$ and an opacity coefficient $\boldsymbol{\gamma}$ that controls visibility over time. This bidirectional design naturally handles the appearance and disappearance of Gaussians without explicitly managing a varying number of Gaussians, thereby facilitating end-to-end training and online updates during inference.
@@
Adaptive Gaussian Fusion via Soft Matching. Directly combining new Gaussians often leads to spatial overlap and redundancy [17]. Traditional optimization-based methods resolve this by rigid one-to-one matching and iterative fusion [47], which are computationally expensive and hard to keep spatial structures, potentially resulting in long-term accumulation of errors.
@@
Inspired by defining the life-cycle of Gaussians [70], we propose an adaptive Gaussian fusion mechanism based on opacity deformation. Each two-frame interval is normalized to $t\in[0,1]$, with $t=0$ and $t=1$ representing the previous and current frames, respectively. Each Gaussian persists across two consecutive frames, with a time-dependent opacity deformation:
@@
$\boldsymbol{\alpha}(t)=\boldsymbol{\alpha}\cdot\frac{\sigma\left(-\boldsymbol{\gamma}_{0}\left(|t-t_{0}|-\boldsymbol{\gamma}_{1}\right)\right)}{\sigma\left(\boldsymbol{\gamma}_{0}\cdot\boldsymbol{\gamma}_{1}\right)},$ (1)
where $\sigma(\cdot)$ denotes sigmoid function, $\boldsymbol{\alpha}$ denotes initial opacity, $t_{0}\in\{0,1\}$ denotes the Gaussian’s creation frame, $\boldsymbol{\gamma}_{0}\in\mathbb{R}^{+}$ and $\boldsymbol{\gamma}_{1}\in[0,1]$ control the transition rate and fade-out timing. This allows overlapped Gaussians from adjacent frames to be implicitly merged through opacity deformation, and the rendering loss can naturally drive the model to find a soft match between them.
Dynamic Deformation Decoding. Given two consecutive frames from a randomly sampled time interval, $I_{t_{1}}$ and $I_{t_{2}}$, we first use the frozen static encoder to extract their 3DGS Embeddings $\mathbf{h}_{t_{1}}$ and $\mathbf{h}_{t_{2}}$. We also extract DINOv2 features [71] $\mathbf{f}_{t_{1}}$ and $\mathbf{f}_{t_{2}}$ from $I_{t_{1}}$ and $I_{t_{2}}$, respectively. To distinguish the embeddings, we add learnable Type Embeddings $\mathbf{T}_{src},\mathbf{T}_{tgt}\in\mathbb{R}^{D}$ to $\mathbf{h}_{t_{1}}$ and $\mathbf{h}_{t_{2}}$, yielding $\hat{\mathbf{h}}_{t_{1}}$ and $\hat{\mathbf{h}}_{t_{2}}$. These are processed by the decoder as follows:

### v2

@@ Abstract
Real-time reconstruction of dynamic 3D scenes from uncalibrated video streams demands robust online methods that recover scene dynamics from sparse observations under strict latency and memory constraints. Yet most dynamic reconstruction methods rely on hours of per-scene optimization under full-sequence access, limiting practical deployment. In this work, we introduce StreamSplat, a fully feed-forward framework that instantly transforms uncalibrated video streams of arbitrary length into dynamic 3D Gaussian Splatting (3DGS) representations in an online manner. It is achieved via three key technical innovations: 1) a probabilistic sampling mechanism that robustly predicts 3D Gaussians from uncalibrated inputs; 2) a bidirectional deformation field that yields reliable associations across frames and mitigates long-term error accumulation; 3) an adaptive Gaussian fusion operation that propagates persistent Gaussians while handling emerging and vanishing ones. Extensive experiments on standard dynamic and static benchmarks demonstrate that StreamSplat achieves state-of-the-art reconstruction quality and dynamic scene modeling. Uniquely, our method supports the online reconstruction of arbitrarily long video streams with a $1200\times$ speedup over optimization-based methods. Our code and models are available at https://streamsplat3d.github.io/.
Figure 1: Given an uncalibrated video stream, StreamSplat instantly reconstructs a dynamic 3D Gaussian scene in an online manner, enabling continuous-time 3D reconstruction, depth estimation, and novel view synthesis.
@@ 1 Introduction
Real-time dynamic 3D reconstruction, a.k.a. 4D reconstruction, from video streams is crucial for numerous applications, e.g., robotics Huang et al. (2023; 2024), augmented/virtual reality (AR/VR) Carmigniani and Furht (2011); Cruz-Neira et al. (1992), and autonomous driving Sun et al. (2020). Specifically, AR/VR systems rely on accurate, continuously updated 3D scenes for immersive experiences, while robots and autonomous vehicles require real-time representations of dynamic environments for safe navigation and responsive interaction.
The prevailing approach to dynamic 3D reconstruction, however, still relies on offline, per-scene optimization Park et al. (2021a; b). Although recent dynamic 3D Gaussian Splatting (3DGS) methods Stearns et al. (2024); Lei et al. (2025) have reduced processing time from days to hours, they continue to follow a multi-step offline pipeline: (i) camera calibration Schonberger and Frahm (2016) and optimization of static 3D Gaussians for each key frame; (ii) optimization of a learnable per-Gaussian deformation field across the entire sequence to capture intermediate dynamics, with the Gaussian set fixed Luiten et al. (2024); Duan et al. (2024); Wang et al. (2024a); and (iii) application of temporal aggregation and fusion (e.g. divide-and-conquer Stearns et al. (2024), k-NN Lei et al. (2025), or interpolation Lee et al. (2024)) to ensure temporal coherence.
Despite recent advances, existing pipelines remain fundamentally offline. They require access to the entire video sequence and hours of iterative, per-scene computation, making them impractical for real-world applications under sparse observations and strict latency constraints. This challenge leads to a crucial research question: Can we match the quality and functionality of offline methods while operating fully online with uncalibrated video streams?
We answer this question affirmatively by introducing StreamSplat, an efficient, scalable, and feed-forward framework for online dynamic 3D reconstruction. Given uncalibrated video streams, StreamSplat directly predicts dynamic 3DGS representations in near real-time. Crucially, it preserves core principles of dynamic 3D reconstruction: (i) maintaining a persistent 3D state via canonical 3D Gaussians predicted from observations, (ii) modeling dynamics through a per-Gaussian deformation field, and (iii) enforcing temporal coherence through streaming aggregation and fusion (Figure 1).
To enable this online reconstruction capability, we propose three technical innovations: 1) Probabilistic position sampling: A mechanism that robustly predicts 3D Gaussians from uncalibrated inputs. This approach captures geometric uncertainty and avoids local minima common in feed-forward models Charatan et al. (2024) (Section 3.1). 2) Bidirectional deformation field: A method that estimates both forward and backward motion between the canonical state and the current observation. This yields reliable associations across frames and mitigates long-term error accumulation (Section 3.2). 3) Adaptive Gaussian fusion: An operation naturally suited for streaming data that performs soft matching to propagate persistent Gaussians while effectively handling emerging and vanishing ones (Section 3.2). In summary, our key contributions are:
@@
  • We introduce StreamSplat, an efficient, scalable, feed-forward framework for online dynamic 3D reconstruction from uncalibrated video streams.
@@
  • We propose three technical innovations: 1) a probabilistic sampling mechanism for 3DGS position prediction, 2) a bidirectional deformation field for robust, efficient dynamic modeling, and 3) an adaptive fusion to maintain a temporally coherent dynamic 3DGS representation.
  • We evaluate our framework on both dynamic (DAVIS Pont-Tuset et al. (2017), YouTube-VOS Xu et al. (2018)) and static (CO3Dv2 Reizenstein et al. (2021), RealEstate10K Zhou et al. (2018)) benchmarks, across video reconstruction, frame interpolation, and novel view synthesis. Our method achieves state-of-the-art performance with a $1200\times$ speedup over optimization-based baselines, while uniquely supporting fully online reconstruction of arbitrarily long streams (Section 4).
@@ 2 Related Works
Dynamic 3D Reconstruction. Early approaches use implicit MLPs optimized per scene Ouyang et al. (2024); Chen et al. (2021); Sitzmann et al. (2020); Tancik et al. (2020), or add time-conditioned deformation fields Pumarola et al. (2021); Park et al. (2021a; b); Li et al. (2022), but still struggle with large, complex motion. Recent methods adopt explicit 3D primitives, particularly dynamic Gaussian Splattings Luiten et al. (2024); Li et al. (2024); Duan et al. (2024); Yuan et al. (2025); Sun et al. (2024); Lee et al. (2024); Shi et al. (2025), for efficient, real-time rendering. However, they still rely on calibrated camera poses and extensive per-scene optimization, making them unsuitable for fully uncalibrated and real-time scenarios.
Feed-Forward Reconstruction. By directly predicting 3D scene representations via neural networks, feed-forward methods have emerged as promising alternatives to optimization-based approaches. Recent SLAM-based Matsuki et al. (2024); Yan et al. (2024); Yugay et al. (2023); Zhu et al. (2024) and scene-coordinate-based methods Wang et al. (2024b); Chen et al. (2025); Jang et al. (2025); Leroy et al. (2024); Xu et al. (2024b); Smart et al. (2024); Yang et al. (2025); Wang et al. (2025a) can estimate camera parameters and 3D structure from uncalibrated inputs but target only static scenes. MonST3R extends to dynamic scene estimations Zhang et al. (2024a), but typically requires pose/geometry post-optimization, and coordinate-based point clouds remain hard to deform Luiten et al. (2024); Duan et al. (2024). On the other hand, feed-forward 3DGS methods achieve fast static reconstruction Wang et al. (2024c); Chen et al. (2024); Zhang et al. (2024b); Shen et al. (2025a); Yi et al. (2024); Xu et al. (2024a); Hong et al. (2024); Zhang et al. (2025); Liu et al. (2024): pixelSplat regresses Gaussians from calibrated pairs Charatan et al. (2024), NoPoSplat handles uncalibrated pairs Ye et al. (2024), and StreamGS enables online static streams Li et al. (2025). However, recent dynamic variants Liang et al. (2024); Yang et al. (2024a); Shen et al. (2025b) still rely on camera calibration and full-sequence access, thus failing to support uncalibrated online streaming. In contrast, StreamSplat operates strictly online on uncalibrated inputs, maintaining temporally consistent dynamic Gaussians via bidirectional deformation and adaptive Gaussian fusion.
@@ 3 Method
In this section, we present StreamSplat, a framework designed to instantly transform uncalibrated online video streams into dynamic 3D Gaussian Splatting (3DGS) representations capable of capturing scene dynamics. Figure 2 provides an overview. We first encode the current frame into static 3D Gaussians in a canonical space (Section 3.1), then predict a bidirectional deformation field between Gaussians encoded from current frame and propagated from the previous frame, and finally fuse them into a unified dynamic representation (Section 3.2). This representation supports rendering at arbitrary times and viewpoints, thereby effectively recovering scene dynamics (Section 3.3).
@@ 3.1 Probabilistic 3D Gaussian Encoding
Canonical 3D Space. To support a unified model for in-the-wild videos with unknown and varying intrinsics (e.g., rectilinear/pinhole and fisheye), we adopt a shared orthographic canonical space following Wang et al. (2023); Sun et al. (2024); Shen et al. (2025b) (details in Appendix A). This choice bypasses per-scene camera calibration: camera motion and perspective effects are absorbed into the Gaussian dynamics and handled by our Dynamic Decoder (Section 3.2). For a detailed discussion on perspective projection, see Appendix D. This canonical space provides a simplified, yet effective foundation for predicting dynamic 3D representations directly from uncalibrated videos.
@@
Structured Static 3D Gaussian Encoding. To overcome the inherent depth ambiguity under orthographic projection Chen et al. (2016) and unstructured nature of 3DGS Chung et al. (2024), we incorporate a pretrained depth estimator Yang et al. (2024b) and predict 3D Gaussian positions in a pixel-aligned manner Charatan et al. (2024).
@@
Given an input frame $I$, we form an RGB-D image (via a pseudo-depth) and partition it into $8\times 8$ patches. A Transformer-based Static Encoder Vaswani et al. (2017) produces 3DGS embeddings $\mathbf{h}$. A lightweight upsampler Liu et al. (2021); Xu et al. (2024c) then yields per-$2\times 2$-patch Gaussian tokens $\hat{\mathbf{E}}$. Each token $\hat{\mathbf{E}}_{i}$ is decoded by linear heads into 3DGS parameters: position offset $\bm{o}_{i}\in[-1,1]^{3}$, rotation $\mathbf{R}_{i}$, scale $\mathbf{S}_{i}$, opacity $\bm{\alpha}_{i}$, and color $\bm{c}_{i}$. The final 3D position is computed via pixel-aligned prediction: $\bm{\mu}_{i}=\left(u_{i}+\bm{o}_{i,0},v_{i}+\bm{o}_{i,1},g(\bm{o}_{i,2})\right),$ where $(u_{i},v_{i})$ is the pixel coordinate, $\bm{o}_{i,2}$ denotes the inverse depth Yang et al. (2024b) for better depth estimation near the camera, and $g(z)=2/(1+z)$ is the depth mapping. Algorithm 1 includes this procedure and detailed architecture is in Figure 9.
Figure 2: Overview of the StreamSplat. Given a pair of frames ($t_{1}=0,t_{n}=1$), we first encode them using the Static Encoder to produce canonical 3D Gaussians (Section 3.1), and then pass the 3DGS Embeddings to the Dynamic Decoder to predict the deformation field (Section 3.2). The predicted dynamic 3D Gaussians can be rendered at arbitrary time $t\in[0,1]$.
@@
Probabilistic Position Sampling. 3D Gaussian Splatting is sensitive to position initialization Kerbl et al. (2023) and prone to local minima Charatan et al. (2024), especially in feed-forward models Charatan et al. (2024); Yi et al. (2024) that make predictions in a single forward pass without iterative refinement. Inspired by Charatan et al. (2024), we predict a truncated normal distribution for each 3D offset $\bm{o}$ rather than regressing it directly: $\bm{o}\sim\mathcal{N}_{[-1,1]}(\bm{\mu}_{p},\bm{\Sigma}_{p})$, where $\bm{\mu}_{p}$ and $\bm{\Sigma}_{p}$ are the predicted mean and covariance. As shown in Section 4.4, this strategy promotes spatial exploration during early training and stabilizes convergence toward optimal positions.
@@ 3.2 Dynamic Deformation Prediction
Online dynamic reconstruction involves large non-rigid motions and topological changes (e.g., emerging/vanishing surfaces). We address this with two modules: (i) a bidirectional deformation field that provides reliable cross-frame associations and naturally handles newly appearing and disappearing content; and (ii) an adaptive Gaussian fusion mechanism that implicitly merges and propagates persistent Gaussians, maintaining long-term temporal coherence. Per-Gaussian deformation is parameterized by a 3D velocity $\mathbf{v}\in[-1,1]^{3}$ and an opacity coefficient $\bm{\gamma}$ that controls visibility over time. Likewise, $\mathbf{v}$ is sampled by the same probabilistic position sampling mechanism (Algorithm 1).
Bidirectional Deformation Field. When a new frame $I_{t}$ arrives, optimization-based methods Duan et al. (2024); Lei et al. (2025) typically instantiate new Gaussians and refine them iteratively. However, this strategy is hard to cast into a feed-forward model: learning a variable number of Gaussians typically requires model selection Akaike (1998) or nonparametric priors like Dirichlet processes Ferguson (1973)), which complicate the end-to-end training.
Instead, we propose to jointly model forward and backward motion between consecutive frames: the forward field deforms previous-frame Gaussians $\mathcal{G}_{t-1}$ to current time $t$, and the backward field deforms current-frame Gaussians $\mathcal{G}_{t}$ back to previous time t1t-1. This symmetric formulation yields robust cross-frame associations and naturally handles emerging and vanishing Gaussians in a unified manner, thereby simplifying prediction and supervision for end-to-end training and online inference.
Figure 3: Our opacity deformation jointly models persistent, emerging, and vanishing Gaussians.
@@
Adaptive Gaussian Fusion via Soft Matching. Directly combining new Gaussians often leads to spatial overlap and redundancy Li et al. (2025). Traditional optimization-based methods resolve this by rigid one-to-one matching and iterative fusion Lee et al. (2024); Stearns et al. (2024), which are computationally expensive and hard to keep spatial structures, potentially resulting in long-term accumulation of errors.
@@
Inspired by defining the life-cycle of Gaussians Zhao et al. (2024), we propose an adaptive Gaussian fusion mechanism based on opacity deformation. Each two-frame interval is normalized to $t\in[0,1]$, with $t=0$ and $t=1$ representing the previous and current frames, respectively. Each Gaussian persists across two consecutive frames, with a time-dependent opacity deformation:
$\bm{\alpha}(t)=\bm{\alpha}\cdot\frac{\sigma\left(-\bm{\gamma}_{0}\left(|t-t_{0}|-\bm{\gamma}_{1}\right)\right)}{\sigma\left(\bm{\gamma}_{0}\cdot\bm{\gamma}_{1}\right)},$ (1)
where $\sigma(\cdot)$ denotes the sigmoid function, $\bm{\alpha}$ denotes initial opacity, $t_{0}\in\{0,1\}$ denotes the Gaussian’s creation frame, $\bm{\gamma}_{0}\in\mathbb{R}^{+}$ and $\bm{\gamma}_{1}\in[0,1]$ controls transition rate and fade-out window. Modulating overlap via time-dependent opacity implicitly fuses the forward and backward Gaussians: the reconstruction loss induces soft match that propagate persistent Gaussians while handling emerging and vanishing ones (Figure 3), thereby maintaining temporal coherence without hard assignments or iterative fusion. In Section 4.2, Figure 4 and Figure 15 further shows that, despite viewpoint/scale changes, motion, occlusions, and blur, our adaptive Gaussian fusion still preserves long-term temporal coherence with persistent Gaussians.
Dynamic Deformation Decoding. Given two consecutive frames $I_{t-1}$ and $I_{t}$, we use the frozen static encoder to obtain 3DGS embeddings $\mathbf{h}_{t-1}$ and $\mathbf{h}_{t}$. A Transformer-based Dynamic Decoder conditioned on DINOv2 features Oquab et al. (2023) produces Deformation Embeddings. The same upsampler yields per-$2{\times}2$-patch deformation tokens $\hat{\mathbf{d}}$, which are concatenated with their paired static tokens $\hat{\mathbf{E}}$ and decoded by a small MLP into per-Gaussian velocity and opacity coefficients. These define forward/backward deformation fields that move and fade Gaussians over time, recovering continuous-time scene evolution. Algorithm 2 includes the procedure; see Appendix B for details.
Input : Embedding $\mathbf{h}_{t_{0}}$, $\mathbf{d}_{t_{0}}$, frame time $t_{0}$
Output : Dynamic 3DGS $\mathcal{G}(t)$
1ex// Static
$(\bm{\mu}_{p},\bm{\Sigma}_{p},\mathbf{S},\mathbf{R},\bm{\alpha}_{0},\bm{c})\leftarrow\textsc{Head}(\mathbf{h}_{t_{0}})$;
$\bm{o}_{0},\bm{o}_{1},\bm{o}_{2}]\sim\mathcal{N}_{[-1,1]}(\bm{\mu}_{p},\bm{\Sigma}_{p})$;
$\bm{\mu}_{0}\leftarrow[u+\bm{o}_{0},v+\bm{o}_{1},g(\bm{o}_{2})]$;
1ex// Dynamic
$(\Delta\bm{\mu}_{p},\Delta\bm{\Sigma}_{p},[\bm{\gamma}_{0},\bm{\gamma}_{1}])\leftarrow\textsc{Head}(\mathbf{h}_{t_{0}},\mathbf{d}_{t_{0}})$;
$\mathbf{v}\sim\mathcal{N}_{[-1,1]}(\Delta\bm{\mu}_{p},\Delta\bm{\Sigma}_{p})$;
1ex// Deformation
$\bm{\mu}(t)\leftarrow\bm{\mu}_{0}+\mathbf{v}\cdot(t-t_{0})$;
@@
$\bm{\alpha}(t)\leftarrow\bm{\alpha}_{0}\cdot\dfrac{\sigma\big(-\bm{\gamma}_{0}\left(|t-t_{0}|-\bm{\gamma}_{1}\right)\big)}{\sigma\left(\bm{\gamma}_{0}\cdot\bm{\gamma}_{1}\right)}$;
$\mathcal{G}(t)\leftarrow\{(\bm{\mu}(t),\bm{\alpha}(t),\mathbf{S},\mathbf{R},\bm{c})\}$;
1exreturn $\mathcal{G}(t)$
Algorithm 1 Dynamic 3DGS Prediction
Input : Frame $I_{k}$, canonical Gaussians $\tilde{\mathcal{G}}(t)$, stored $\mathbf{h}_{k-1}$, camera pose $\pi$
Output : Updated $\tilde{\mathcal{G}}(t)$, stored $\mathbf{h}_{k}$, rendered frames $\hat{I}_{t_{k-1}\rightarrow t_{k}}$
// Prediction
$\mathbf{h}_{k}\leftarrow\textsc{Encoder}(I_{k})$;
$(\mathbf{d}^{+}_{k-1},\mathbf{d}^{-}_{k})\leftarrow\textsc{Decoder}(\mathbf{h}_{k-1},\mathbf{h}_{k})$;
$\mathcal{G}_{k-1}^{+}(t)\leftarrow\textsc{Pred}(\mathbf{h}_{k-1},\mathbf{d}^{+}_{k-1},t_{k-1})$;
$\mathcal{G}_{k}^{-}(t)\leftarrow\textsc{Pred}(\mathbf{h}_{k},\mathbf{d}^{-}_{k},t_{k})$;
// Alg. 1
// Update Deformation & Fusion
$\tilde{\mathcal{G}}(t)\leftarrow\textsc{Update}\big(\tilde{\mathcal{G}}(t),\mathcal{G}_{k-1}^{+}(t)\big)\cup\mathcal{G}_{k}^{-}(t)$;
$\hat{I}_{t_{k-1}\rightarrow t_{k}}\leftarrow\textsc{Rasterize}(\tilde{\mathcal{G}}(t),\pi)$;